OpenStudy (anonymous):

Write the equation for the horizontal line passing through (4, -6).

OpenStudy (lgbasallote):

horizontal line means all your x coordinates are the same so...your equation yould be y = your y coordinate

OpenStudy (lgbasallote):

your equation would*

OpenStudy (anonymous):

x=4 y=-6

OpenStudy (zepp):



OpenStudy (zepp):

If it's an horizontal line, the slope is 0, your function would be like f(x) = 0x + yintercept We know that the y intercept is when x=0, in this question, when x=0, y = -6 So the y intercept is -6 f(x) = 0x -6 f(x) = -6

OpenStudy (anonymous):

these are the answers i can get A.y = 4 B.y = -6 C.y = 4/6 D.y = -6/4

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (zepp):

The answer is right there, if you read my explanations.

OpenStudy (anonymous):

-6